The speed c of ocean wave is proportional to the acceleration due to gravity g, wavelength λ, and density ρ of the wave such that c = kgxλyρz,  where k is a dimensionless constant. What is the correct equation for the speed of the ocean?


1
Expert's answer
2021-01-13T11:36:44-0500

Write the equation:


"c=kg^x\lambda^y\rho^z."

Write the units instead of quantities:


"[\text{m/s}]=1\cdot[\text{m/s}^2]^x[\text{m}]^y[\text{kg/m}^3]^z,\\
[\text{m/s}]=[\text{m}^x/\text{}{s}^{2x}][\text{m}^y][\text{kg}^z/\text{m}^{3z}]."

We see that the left part (m/s) does not contain kilograms, thus, "z=0".

Since we have seconds in the power of 1, "x=1\/2".

Since in the left part we have meters in the power of one and in the right part we have "m^{1\/2}", the product with "m^y" must result in "m^1", or


"1/2+y=1,\\
y=1/2."

Thus, the equation is


"c=k\sqrt{g\lambda}."

Indeed, the equation used to calculate wave velocity is


"c=\frac{1}{\sqrt{2\pi}}\cdot\sqrt{g\lambda}."